OpenStudy (anonymous):

Solve for h. g = 3h - 4i

OpenStudy (anonymous):

where are you stuck at?

OpenStudy (anonymous):

@DemolisionWolf i have no idea how to do this dude, im doing online classes and practically have to turor myself

OpenStudy (anonymous):

tutor ^

OpenStudy (anonymous):

g = 3h - 4i add 4i to both sides g+4i = 3h - 4i + 4i g+4i = 3h now divide both sides by 3 (g+4i)/3 = (3h)/3 (g+4i)/3 = h